# Hybrid Photon-magnon Systems: Exploring the Purcell Effect

### v1

@@ III Results and Discussions
where $\tilde{\omega}_{c}^{\prime}=\tilde{\omega}_{c}-\lambda_{c}$, $\tilde{\omega}_{m}^{\prime}=\tilde{\omega}_{m}-\lambda_{m}$, and $g^{\prime}=g-i\sqrt{\lambda_{c}\lambda_{m}}$.
The eigenvalues of the coupling matrix can be expressed as:
$\tilde{\omega}_{\pm}=\frac{1}{2}\left[\tilde{\omega}_{c}^{\prime}+\tilde{\omega}_{m}^{\prime}\pm\sqrt{\left(\tilde{\omega}_{c}^{\prime}-\tilde{\omega}_{m}^{\prime}\right)^{2}+4g^{\prime 2}}\right].$ (5)
The transmission coefficient ($|S_{21}|$ spectra) for the PMC system, derived from Eq. 1 using input-output theoryRao et al. (2020); Shrivastava et al. (2024), is given by:
@@
$|S_{21}|=\frac{2\alpha\gamma_{c}\omega_{m}+2\beta\gamma_{m}\omega_{c}-4ig\sqrt{\gamma_{c}\gamma_{m}}-2i\gamma_{c}(\omega-\omega_{m})-2i\gamma_{m}(\omega-\omega_{c})}{\left[(ig+\sqrt{\gamma_{c}\gamma_{m}})^{2}+(i\beta\omega_{c}+i\gamma_{c}+\omega-\omega_{c})(i\alpha\omega_{m}+i\gamma_{m}+\omega-\omega_{m})\right]}.$ (6)
@@
Here, $\gamma_{c,m}=2\pi\lambda_{c,m}^{2}$, representing the extrinsic damping rates of HRR and YIG, respectively.
Using Eq. 6, we plotted the $|S_{21}|$ power as a function of frequency for different magnon damping values, as shown in Fig. 3(a). The variation in $|S_{21}|$ power with magnon damping closely aligns with our simulation results depicted in Fig. 2(a), demonstrating excellent agreement between the theoretical model and simulation. We also calculated the photon mode linewidth ($K_{c}/2\pi$) corresponding to magnon damping using our quantum model. From the calculated parameters ($K_{c}/2\pi$), ($g/2\pi$), and ($K_{m}/2\pi$) as listed in Table 1, we found that the Purcell condition, $((K_{m}-K_{c})/2<g\leq K_{m})$, is satisfied only for $\alpha\geq 2.1\times 10^{-2}$. The corresponding values of $K_{c}/2\pi$ are 39 MHz and 45.5 MHz, while $g/2\pi$ are 76.03 MHz and 62.6 MHz, respectively.
This result confirms the occurrence of the Purcell effect at higher magnon damping, further validating the accuracy and consistency of our theoretical model.
Figure 3: (a) $|S21|$ power of photon-magnon coupled modes on f-H plane, calculated using a quantum model for different values of magnon damping. The variation of (b) coupling strength $(g/2\pi)$, and (c) the damping constant of the photon mode $(\beta)$, both as functions of magnon damping $(\alpha)$. (e) The $|S21|$ transmission Spectra as a function of time, are calculated using the FFT of simulation data; and (f) The $|S21|$ spectra as a function of frequency, for two different bias magnetic fields ($H_{dc1}=900Oe$, and $H_{dc2}=1210Oe$) applied along the Y-direction, shown only for the highest magnon damping $(\alpha=2.8\times 10^{-2})$.
To further investigate the variation of coupling strength with magnon damping through quantum calculations, we calculated the coupling strength using Eq. (5) and presented the results as blue stars in Fig. 3(b). The analysis shows that the coupling strength decreases gradually for low magnon damping, reaching 116.61 MHz, but exhibits a pronounced reduction at higher damping, dropping to 62.6 MHz for $\alpha=2.8\times 10^{-2}$. The photon damping rate, calculated using Eq. (5) (see Table 1) and shown in Fig. 3(c), increases significantly with higher magnon damping and exhibits enhanced spontaneous emission only for $\alpha\geq 2.1\times 10^{-2}$.
@@ IV Conclusion
In summary, through numerical simulations, we systematically analyze the role of magnon damping in tailoring the coupled mode dynamics of a photon-magnon hybrid system. By varying the magnon damping, we identify the conditions for the emergence of the Purcell effect, characterized by reduced coupling strength and increased photon dissipation. Our quantum theoretical model findings not only validate the occurrence of the Purcell effect but also reveal new insights into the tunability of hybrid systems by manipulating dissipation channels. This study expands the scope of cavity magnonics, offering a novel pathway to dynamically control hybrid quantum systems for applications in quantum information science and beyond.

### v2

@@ III.2 Theoretical Model and Input-Output Theory
where $\tilde{\omega}_{c}^{\prime}=\tilde{\omega}_{c}-i\gamma_{c}$, $\tilde{\omega}_{m}^{\prime}=\tilde{\omega}_{m}-i\gamma_{m}$, and $g^{\prime}=g-i\sqrt{\gamma_{c}\gamma_{m}}$. The eigenvalues of the coupling matrix can be expressed as:
$\tilde{\omega}_{\pm}=\frac{1}{2}\left[\tilde{\omega}_{c}^{\prime}+\tilde{\omega}_{m}^{\prime}\pm\sqrt{\left(\tilde{\omega}_{c}^{\prime}-\tilde{\omega}_{m}^{\prime}\right)^{2}+4g^{\prime 2}}\right]$ (9a)
[…]
The frequency gap, $\Delta$ between the two modes at the anti-crossing center $(\omega_{c}=\omega_{m})$ is given by $\Delta=(\tilde{\omega}_{+}-\tilde{\omega}_{-})/2\pi$. Using Equation (9a), $\Delta$ can be expressed as:
@@
$\Delta=\frac{1}{2\pi}\left[\sqrt{4\left(g^{\prime}\right)^{2}-\left[\omega_{c}\left(\beta-\alpha\right)+\left(\gamma_{c}-\gamma_{m}\right)\right]^{2}}\right]$ (9b)
To calculate the transmission coefficient ($|S_{21}|$) of the PMC system, the steady-state equations in the frequency domain are obtained by applying a Fourier transform to Eq. (7), which gives:
$i(\omega-\tilde{\omega_{c}})\hat{c}(\omega)-i\sqrt{\gamma_{c}}\,\hat{p}_{in}(\omega)-\gamma_{c}\,\hat{c}(\omega)-\sqrt{\gamma_{c}\gamma_{m}}\,\hat{m}(\omega)$
$-ig\hat{m}(\omega)=0$ (10a)
$i(\omega-\tilde{\omega_{m}})\hat{m}(\omega)-i\sqrt{\gamma_{m}}\,\hat{p}_{in}(\omega)-\gamma_{m}\,\hat{m}(\omega)-\sqrt{\gamma_{c}\gamma_{m}}\,\hat{c}(\omega)$
$-ig\hat{c}(\omega)=0$ (10b)
From Eq. (4a), for ${t_{1}}>t$, the traveling photon $\hat{p}_{k}$ after its interaction with the two modes is expressed as:
$\hat{p}_{k}(t)=e^{-i\omega_{k}(t-t_{1})}\hat{p}_{k}(t_{1})-\int_{t}^{t_{1}}i\big[\lambda_{c}\hat{c}+\lambda_{m}\hat{m}\big]e^{-i\omega_{k}(t-t^{\prime})}dt^{\prime}$ (11a)
The field operator corresponding to the output port is defined as:
$\hat{p}_{out}(t)=\frac{1}{\sqrt{2\pi}}\int e^{-i\omega_{k}(t-t_{1})}\hat{p}_{k}(t_{1})\,dk$ (11b)
By substituting Eq. (11) into Eq. (5), the time-retarded Heisenberg–Langevin equations for the coupled system can be obtained in both the time and frequency domains.
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}=-i\begin{bmatrix}\tilde{\omega}_{c}+i\gamma_{c}&g+i\sqrt{\gamma_{c}\gamma_{m}}\\ g+i\sqrt{\gamma_{c}\gamma_{m}}&\tilde{\omega}_{m}+i\gamma_{m}\end{bmatrix}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}-i\begin{bmatrix}\sqrt{\gamma_{c}}\\ \sqrt{\gamma_{m}}\end{bmatrix}\hat{p}_{\text{out}}(t).$ (12)
Applying a Fourier transform to Eq. (12) yields:
$i(\omega-\tilde{\omega_{c}})\hat{c}(\omega)-i\sqrt{\gamma_{c}}\,\hat{p}_{out}(\omega)+\gamma_{c}\,\hat{c}(\omega)+\sqrt{\gamma_{c}\gamma_{m}}\,\hat{m}(\omega)$
$-ig\hat{m}(\omega)=0$ (13a)
$i(\omega-\tilde{\omega_{m}})\hat{m}(\omega)-i\sqrt{\gamma_{m}}\,\hat{p}_{out}(\omega)+\gamma_{m}\,\hat{m}(\omega)+\sqrt{\gamma_{c}\gamma_{m}}\,\hat{c}(\omega)$
$-ig\hat{c}(\omega)=0$ (13b)
By solving Eqs. (10) and (13), the relation between the input and output fields of this coupled system can be obtained as:
$\hat{p}_{out}=\hat{p}_{in}-2i\sqrt{\gamma_{c}}\,\hat{c}(\omega)-2i\sqrt{\gamma_{m}}\,\hat{m}(\omega)$ (14)
Solving Eqs. 10(a) and (b) provides the field operators expressed in terms of the input field as:
$\hat{c}(\omega)=\frac{\hat{p}_{in}\left[g\sqrt{\gamma_{m}}+i\alpha\sqrt{k_{c}}\,\omega_{m}+\sqrt{k_{c}}\,\omega-\sqrt{k_{c}}\,\omega_{m}\right]}{\left[(ig+\sqrt{\gamma_{c}\gamma_{m}})^{2}+(i\beta\omega_{c}+i\gamma_{c}+\omega-\omega_{c})(i\alpha\omega_{m}+i\gamma_{m}+\omega-\omega_{m})\right]}$ (15a)
$\hat{m}(\omega)=\frac{\hat{p}_{in}\left[g\sqrt{\gamma_{c}}+i\beta\sqrt{k_{m}}\,\omega_{c}+\sqrt{k_{m}}\,\omega-\sqrt{k_{m}}\,\omega_{c}\right]}{\left[(ig+\sqrt{\gamma_{c}\gamma_{m}})^{2}+(i\beta\omega_{c}+i\gamma_{c}+\omega-\omega_{c})(i\alpha\omega_{m}+i\gamma_{m}+\omega-\omega_{m})\right]}$ (15b)
Therefore, the transmission coefficient of the coupled system can be expressed as:
$|S_{21}|=\frac{\hat{p_{out}}}{\hat{p}_{in}}-1$ (16a)
$\|S_{21}|=\frac{2\alpha\gamma_{c}\omega_{m}+2\beta\gamma_{m}\omega_{c}-4ig\sqrt{\gamma_{c}\gamma_{m}}-2i\gamma_{c}(\omega-\omega_{m})-2i\gamma_{m}(\omega-\omega_{c})}{\left[(ig+\sqrt{\gamma_{c}\gamma_{m}})^{2}+(i\beta\omega_{c}+i\gamma_{c}+\omega-\omega_{c})(i\alpha\omega_{m}+i\gamma_{m}+\omega-\omega_{m})\right]}$ (16b)
Figure 3: (a) $|S_{21}|$ transmission coefficient of photon-magnon coupled modes on fHf-H plane, calculated using a quantum model for different values of magnon damping. The variation of (b) coupling strength $(g/2\pi)$, and (c) the damping constant of the photon mode $(\beta)$, both as functions of magnon damping $(\alpha)$. (d) The $|S_{21}|$ transmission Spectra as a function of time, are calculated using the FFT of simulation data; and (e) The $|S_{21}|$ spectra as a function of frequency, for two different bias magnetic fields ($H_{dc1}=900Oe$, and $H_{dc2}=1210Oe$) applied along the Y-direction, shown only for the highest magnon damping $(\alpha=2.8\times 10^{-2})$.
@@
III.3 Numerical verification of Purcell Condition
To further validate the analytical framework, we examine the numerical signatures of the Purcell effect by evaluating the transmission spectra using Eq. (16). Figure 3(a) shows the calculated $|S_{21}|$ transmission amplitude coefficient as a function of frequency for different magnon damping values which closely follows the simulated results in Fig. 2(a), demonstrating a good agreement between the theoretical model and the simulations. From the calculated photon mode linewidth $K_{c}/2\pi$, coupling strength $g/2\pi$, and magnon linewidth $K_{m}/2\pi$, summarized in Table 1, we find that the Purcell condition, $((K_{m}-K_{c})/2<g\leq K_{m})$, is satisfied only for $\alpha\geq 2.1\times 10^{-2}$. The corresponding values of $K_{c}/2\pi$ are 39 MHz and 45.5 MHz, while $g/2\pi$ are 76.03 MHz and 62.6 MHz, respectively. These values confirm that the Purcell effect emerges at higher magnon damping, thereby validating both the accuracy and predictive power of our theoretical model.
To further examine the influence of magnon damping on coupling dynamics, we calculated the coupling strength using Eq. (9), with the results shown in Fig. 3(b). The analysis shows that the coupling strength decreases gradually for low magnon damping, reaching 116.61 MHz, but exhibits a pronounced reduction at higher damping, dropping to 62.6 MHz for $\alpha=2.8\times 10^{-2}$. The photon damping rate, calculated using Eq. (9) (see Table 1) and shown in Fig. 3(c), increases significantly with higher magnon damping and exhibits enhanced spontaneous emission only for $\alpha\geq 2.1\times 10^{-2}$.
Figure 4: (a) Extracted values of $\beta$ as a function of spin density, showing systematic enhancement of magnon damping with increasing spin reservoir, consistent with the Purcell effect. (b) Calculated coupling strength $g$ versus total spin number NN, exhibiting the expected $g=g_{0}\sqrt{N}$ scaling characteristic of collective coherent interactions.
@@ III.3 Numerical verification of Purcell Condition
III.4 Control of Purcell Effect
To further explore the controllability of the Purcell effect in planar PMC systems, we examined its dependence on spin density by tuning the spin volume through controlled variation of the YIG film thickness. The corresponding color maps of the transmission spectra for different spin volumes are presented in Supplementary Fig. S3(a), with theoretical simulations shown in Fig. S3(b). Using the quantum model and Eq. (9b), the values of $\beta$ were extracted as a function of spin density [Fig. 4(a)], which clearly demonstrate that the damping of the HRR increases systematically with larger spin reservoirs. This behavior signifies that the Purcell effect manifested as enhanced photon dissipation becomes more dominant when the collective spin ensemble is larger, thereby accelerating photon leakage into the magnonic channel. In parallel, the coupling strength obtained from simulations, shown in Fig. 4(b), follows the expected square-root law $g=g_{0}\sqrt{N}$, with $N$ being the total number of spins. This $g=g_{0}\sqrt{N}$ scaling is a hallmark of collective coherent interactions, confirming that the enhancement of the Purcell effect originates from cooperative spin ensemble dynamics rather than single-spin processes. These findings highlight that both magnon damping and spin density serve as tunable parameters for dynamically controlling the Purcell effect, thus offering a multidimensional design space for optimizing photon–magnon hybrid platforms in quantum technologies.
While the above analysis clarified the influence of magnon damping and spin density, a more comprehensive framework is needed to delineate the parameter regimes governing the transition from coherent coupling to the Purcell regime. To this end, we mapped the interdependence of $\alpha,\beta$, and $g$ in a unified parameter space. Since the coupling under investigation is coherent, we focus only on the real part of $\Delta$ to capture the effective mode splitting. Figure 5, obtained from Eq. (9b) with $\omega_{c}=5.33$ GHz, presents the phase diagram where the color gradient of $Re(\Delta)$ visualizes the effective coupling strength: red–yellow regions correspond to strong coupling, whereas blue regions indicate weak interaction. This representation clearly illustrates how increasing magnon damping systematically drives the system from strong coherent coupling toward the Purcell regime. As expected, coupling strength decreases monotonically with higher $\alpha$, with deep-blue regions denoting the weakest coupling at large $\alpha$ and red regions highlighting the strongest interaction at low $\alpha$. The black star markers, representing $\beta$ values calculated for different $\alpha$, align within the red–orange volume, confirming their placement in the cavity-enhanced decay regime. Although these points do not fully satisfy the Purcell condition, their systematic shift toward higher $\beta$ with increasing $\alpha$ outlines a clear transition pathway into the Purcell-dominated regime. Furthermore, the shaded region in Fig. 5 marks the parameter space where the Purcell condition is met, showing that the two highest damping cases lie squarely within this region, while the shaded region in Fig. S4 of the Supplementary provides a magnified view of the same parameter space where the Purcell condition is satisfied. Together, these results provide a comprehensive visualization of how $\alpha,\beta$, and $g$ dictate the system’s operational regime, establishing a clear framework for engineering transitions between coherent coupling and Purcell-enhanced dissipation in planar PMC systems.
Our results demonstrate that achieving the Purcell effect in a photon-magnon coupled hybrid system requires the use of lossy magnetic materials. The damping of YIG can be tuned over a wide range $10^{-1}$ to $10^{-5}$ through various techniques, including growth optimization, rare-earth metal doping, spin pumping, nonlocal spin currents, and substrate engineering. Growth optimization, achieved by modifying deposition and annealing conditions or using different substrates, enables intrinsic control over dampingBhoi et al. (2018). Rare-earth doping introduces impurity scattering, influencing relaxation dynamicsGurjar et al. (2021), while spin pumping provides an external approach where a spin-polarized current, injected via a nanocontact, dynamically modulates the dampingSun et al. (2013); Heinrich et al. (2011); Navabi et al. (2019). Additionally, nonlocal spin currents and substrate orientation adjustments fine-tune Gilbert damping, as demonstrated in prior studiesGurjar et al. (2021); Haidar et al. (2016).
Figure 5: Analytically derived three-dimensional phase diagram of the real part of $\Delta$ depicting various forms of anti-crossing dispersion within the $\left(\alpha-\beta-g\right)$ parameter space. The color gradient represents the variation of the coupling strength $\Delta$, while the green-shaded region highlights the Purcell regime. Theoretical $\beta$ values are marked with solid black star marker with white outline.
Temperature is another critical factor influencing ferromagnetic resonance linewidth and magnon lifetime. While our analysis has been based on room-temperature parameters most relevant for practical on-chip applications magnon damping exhibits strong temperature dependence. For instance, Jermain et al. reported a substantial increase in linewidth in thin-film YIG at low temperatures due to impurity relaxationJermain et al. (2017), whereas Kosen et al. demonstrated exceptionally low linewidths (1.4 MHz) in high-quality YIG at millikelvin temperatures, dominated by coupling to two-level fluctuators Kosen et al. (2019). These studies underscore that although room-temperature operation ensures technological relevance, cryogenic operation can further enhance magnon coherence and open additional opportunities for quantum devices.
A comparison between our obtained coupling strengths and previously reported values for millimetre-scale YIG magnets is summarized as follows. Harder et. al. observed a coupling strength of 31.5 MHz for a YIG sphere in a 3D cavityHarder et al. (2016), while Hyde et al. reported 30.3 MHz for a similar YIG sphere/3D cavity systemHyde et al. (2016). Kaur et al. demonstrated a larger value of 65 MHz for a YIG sphere coupled to a split-ring resonator on a striplineKaur et al. (2016), and Zhang et al. obtained 32.85 MHz for a YIG cylinder/SRR stripline systemZhang, Song, and Chai (2017). In comparison, the coupling strengths achieved in our planar YIG/HRR hybrid structure are tunable between 62.6 MHz and 126.6 MHz, which are significantly higher than the typical values reported for millimetre-scale spherical or cylindrical YIG resonators. These results indicate that while spherical YIG resonators benefit from large spin ensembles and mode volumes, carefully engineered planar geometries such as the present YIG/HRR design not only enable stronger and tunable coupling but also offer distinct advantages in terms of compatibility with on-chip integration.
Experimental studies have validated many of these strategies. For example, Bhoi et al.Bhoi et al. (2018) demonstrated strain-induced damping modulation in nanometer-thick YIG films grown by pulsed laser deposition, while others achieved damping control via nonlocal spin currentsHaidar et al. (2016) and substrate orientation modificationsGurjar et al. (2021). Notably, Zhang et al.Zhang et al. (2014) observed the Purcell effect in a 3D microwave cavity, where enhanced cavity photon decay was induced by coupling with the lossy magnon mode of a YIG sphere. In that work, losses were deliberately introduced by attaching iron filings to the sphere, thereby increasing scattering and absorption. This experimental evidence strongly supports our findings that the Purcell effect in hybrid photon-magnon systems can be realized through multiple strategies for engineering lossy magnetic materials.
@@ IV Conclusion
In summary, this study establishes a theoretical framework for realizing the Purcell effect in photon–magnon coupled systems, identifying magnon damping as a central tunable parameter for controlling spontaneous emission and photon dissipation. Through numerical simulations and parameter-space mapping, we show how damping, photon loss, and coupling strength collectively determine the transition between coherent coupling and the Purcell regime, while spin-density scaling confirms the cooperative nature of the effect. The feasibility of tuning YIG damping via growth optimization, rare-earth doping, spin pumping, and substrate engineering underscores the experimental practicality of our approach. By integrating magnetic control with Purcell engineering in a planar photon–magnon system, we demonstrate tunable collective light–matter interactions, revealing transitions between strong coupling and the Purcell regime and providing design principles for scalable, chip-compatible hybrid architectures.